# Plasmon Enhanced Spectroscopy and Photocatalysis

### v1

@@ Abstract
This study examines the Raman scattering and charge transfer properties of molecules adsorbed on the surface of a tetrahedral Au nanoparticle based on the time-dependent density functional tight-bonding (TD-DFTB) method.
@@
Surface-enhanced Raman scattering (SERS) is a widely used phenomenon among various research communities where the scattering intensity of molecules can be increased by several orders of magnitude when they are adsorbed on the surface of plasmonic nanoparticles.SERS-VanDuyne-1997 (1, 2, 3)
The enhancement factors are as large as and therefore SERS can be studied on the single-molecule level.SERS-Emory-1997 (4, 5, 6, 7, 8)
The unexpected enhancement in the Raman signal from pyridine adsorbed on an electrochemically roughened silver electrode was the first evidence of SERS,SERS-McQuillan-1974 (9) and this was attributed to the extensive surface coverage of molecules on the rough nanostructured surface, but later it was identified as a plasmon-based phenomenon.SERS-VanDuyne-1997 (1, 2)
SERS has been successfully applied to several fields including biosensingBiosensing-VanDuyne-2003 (10, 11), drug analysis,Drug-Schatz-2009 (12) and electrochemistryElectrochemical-Zen-2015 (13, 14).
@@
Localized surface plasmon resonance (LSPR) excitation of the conduction electrons in metal nanoparticles of Au, Ag, and Cu with dimensions of 5-200 nanometers generates an enhanced electromagnetic field at the interface of the particle.LSPR-VanDuyne-2007 (15)
This enhanced field can modulate several optical processes at the surface including photocurrent,Photocurrent-Schatz-2009 (16) luminescence,Exciton-Achermann-2010 (17) and Raman scattering.SERS-VanDuyne-1997 (1, 2, 3)
Raman scattering scales with the local field as , where the numerator is the localized electromagnetic field and the denominator is the incident field.SERS-VanDuyne-2008 (3)
SERS enhancement factors due to electromagnetic effects can reach values up to .SERS-Schatz-2008 (6)
@@
Charge transfer (CT) excitations at optical frequency can also contribute to enhancements through a resonance Raman mechanism.Electronic-Schatz-2008 (18)
Enhancements due to static chemical effects can be up to a factor of 3-20 as demonstrated theoretically for pyridine adsorbed on the surface of a small tetrahedral Au particle.TDDFT-Schatz-2006 (19). Resonant CT can lead to larger enhancements, but usually the resonant CT and LSPR frequencies are sufficiently different, or the resonance CT transition is sufficiently broad, that one only has significant SERS enhancement from plasmon excitation and static CT.
@@
Previous theoretical works have addressed SERS for pyridine adsorbed on the surface of Ag and Au particles with enhancement factors up to based on the time-dependent density functional theory (TD-DFT).TDDFT-Schatz-2006 (19, 19)
@@
Semi-empirical methods based on INDO have been demonstrated successfully for the efficient simulation of SERS for Ag-pyridine complexes.HyperRaman-Schatz-1992 (20, 21)
As an alternative to the INDO calculations, the density functional tight-binding (DFTB) method may also provide efficient solutions for larger complexes.DFTB-Aradi-2020 (22)
@@
Two recent papers provide details of the DFTB method.DFTB-Aradi-2020 (22, 23)
@@
Previous works have demonstrated the success of the DFTB method in various applications including biomolecules,DFTB-Bio-Wong-2019 (24) metals,DFTB-Metal-2019 (25, 26, 27) semiconductors,DFTB-SemiConductor (28) plasmonic photocatalysis,PDC-Sanchez-2022 (29, 30, 31), etc.
@@
The real-time time-dependent density functional tight-binding (RT-TD-DFTB) method allows us to study plasmon-driven charge and energy transfer processes at the metal-molecule interface by simulating laser-induced electron dynamics.RTDFTB-Sanchez-2020 (23, 29, 30)
@@
This includes several processes like molecular bond dissociation and formation,PDC-Schatz-2023 (30, 31) charge transfer (CT),Hot-Carrier-Erhart-2022 (32) energy transfer,PDC-Sanchez-2022 (29) spectral shift,SERS-PCBM-Schatz-2019 (33) etc.
In the following, we study CT processes between Au and a fullerene (C) after exciting plasmons at 2.5 eV with an external laser pulse.
This work was inspired by a recent study from the Van Duyne group, where the fullerene-based molecule [6,6]-phenyl-C-butyric acid methyl ester (PCBM) was found to form negative ions in the presence of gold dimers when irradiated at 532 nm as revealed by SERS measurements.
@@
The paper is divided into two parts, in the first part, we explore Raman scattering for Au-pyridine complexes and the second part contains plasmon-driven CT in a Au-C complex.
@@ 2 Computational Details
To elucidate the accuracy of this method for predicting Raman spectra, we compare TD-DFTB results with conventional results obtained with the B3LYP/TZVP level of DFT theory.
@@
To obtain a smooth Raman spectrum as a function of Raman shifts, the scattering activity is broadened by a Lorentzian function with a cm width.
In real-time TD-DFTB (RT-TD-DFTB) the one-body reduced density matrix is propagated according to the Liouville-von Neumann equation in a nonorthogonal basisRTDFTB-Sanchez-2020 (23)
@@
with field amplitude , center of the pulse , pulse duration , and frequency . For calculating spectra, the density matrix evolves in time due to purely electronic dynamics where the nuclei are “clamped” in space. This treatment has been extensively used to study optical processes of material and molecules induced by external time-dependent electric fields.RTDFTB-Sanchez-2016 (34, 35, 23) The Mulliken population is calculated as
@@
with an initial permanent dipole .
@@ 3 Results and Discussion
Au-Pyridine Complexes:
We use a tetrahedral Au nanoparticle for all of these studies, which is a similar structure (but larger) to what has been considered in earlier SERS studies.TDDFT-Schatz-2006 (19, 19) This allows two distinct binding sites for an adsorbing pyridine molecule, either on one of the tips (V-complex) or in the middle of the triangle faces (S-complex).
@@
We optimize all geometries with a maximum force tolerance of 10 a.u. for the movement of atoms.
We attach pyridine on the tip (for V-complex) and surface (for S-complex) gold atoms of a previously optimized Au and further optimize, freezing the atoms of the particle.
@@
A smaller distance between the particle and molecule was also observed in a Au-pyridine V-complex with TD-DFT calculations.TDDFT-Schatz-2006 (19)
@@
These spectral features are consistent with previous studies based on INDO/SCI and TD-DFT methods.INDO-Schatz-2017 (21)
@@
Absorption spectra for Au and Au-pyridine complexes are shown in Figure 2(b).
The spectrum for the gold particle shows multiple peaks that are comparable to the spectrum of a tetrahedral Au investigated using the TD-DFT method.TDDFT-Schatz-2006 (19)
@@
However, the overall spectral intensity is increased with the particle size due to the large density of states available in Au.
@@
We note that although a gold particle involving 120 atoms is large compared to the previous study with Au, it is still too small sustain significant LSPR excitation.
@@
The spectrum for the S-complex is very similar to the isolated gold particle spectrum, whereas for the V-complex the absorption intensity is slightly increased due to CT excitations as is consistent with previous TD-DFT results.TDDFT-Schatz-2006 (19)
@@
We have used DFTB to compute the SERS enhancement for pyridine using low frequency (nominally static) calculations to determine the chemical enhancement factor, and with calculations at the plasmon energy of 2.5 eV for Au-pyridine complexes to determine the electromagnetic enhancement factor.
We find that resonant CT excitations for the Au-pyridine complexes involve small transition strengths that overlap with the dominant -band excitations and therefore can have negligible contributions to SERS.
@@
The Raman active modes from the TD-DFT calculation include a pair of ring deformation modes at 620 and 672 cm, a pair of ring breathing modes at 997 and 1039 cm, a symmetric C-H wag mode at 1242 cm, and a pair of ring stretching modes at 1604 and 1605 cm.
The two noticeable discrepancies between spectra are (i) unphysical modes that appear in the TD-DFTB results near 794 cm and (ii) for the lower energy ring breathing mode (that appears at 941 cm in the TD-DFTB result) the intensity is significantly suppressed compared to the others.
As will be discussed later the unphysical modes near 794 cm are not enhanced in SERS, however, the lower energy ring breathing mode can be significantly enhanced even though it is vanishingly small for the isolated gas phase molecule (TD-DFTB result).
The normal Raman scattering spectra of pyridine adsorbed on the surface of Au are shown in Figure 4.
@@
These enhancement factors are comparable with the reported values (a factor 3 for the S-complex and 20 for the V-complex) for a Au-pyridine complex based on earlier TD-DFT calculations.TDDFT-Schatz-2006 (19)
@@
Interestingly the lower energy ring breathing mode (944 cm) that has a small intensity for the isolated molecule (Fig.3) is now enhanced by a factor of 15 for the V-complex, while the higher energy ring breathing mode (1056 cm) shows a weak enhancement.
@@
Again, this trend in enhancements is consistent with the TD-DFT calculation as observed for Au-pyridine complexes.TDDFT-Schatz-2006 (19).
Further, the unphysical modes that appear in the isolated molecule near 794 cm have not been enhanced for either complex.
Another mode with much higher enhancements is near 1459 cm. This peak was also seen in the TD-DFT work, and it appears in experimental SERS measurements on gold but not silver.TDDFT-Schatz-2006 (19)
The mode frequencies show blue-shifts for both complexes as observed in the TD-DFT calculations, however, the shifts are relatively smaller compared to the TD-DFT results.TDDFT-Schatz-2006 (19)
@@
Similar enhancement factors were also observed for Au-pyridine complexes with the TD-DFT method.TDDFT-Schatz-2006 (19) For the V-complex, the 941 and 1635 cm modes are enhanced significantly, by a factor of and , respectively, while for the S-complex multiple modes are enhanced, with the maximum enhancement for 1636 cm mode.
Further, a new Raman active mode appears in the S-complex at 1459 cm.
@@
AuC Complex:
To generate a suitable geometry to study plasmon-driven charge transfer, we place a fullerene (C) molecule on one of the triangular faces (as in the S-complex) of Au (previously optimized) and then the geometry is optimized freezing atoms of the particle.
The optimized geometry of AuC is shown in Figure 6.
@@
Absorption spectra for C and AuC are displayed in Figures 2 (a) and (b), respectively.
@@
These calculated peaks well represent the experimental peaks found at 3.8, 4.9, and 6 eV, respectively for the gas phase C molecule.C60-Shobatake-1996 (36)
The spectrum of the AuC complex is dominated by the gold particle with two major peaks near 3.27 and 4.3 eV.
@@
Previous studies demonstrated a shift in LSPR peaks due to the interaction between plasmon and molecular resonances.Plasmon-Molecule-VanDuyne-2006 (37, 38)
@@
A linearly polarized (polarized along the Z-axis) laser pulse of the form given in Eq.2 is used to drive the system. Both molecule and particle couple with the external field. We use a driving field amplitude , frequency , and pulse duration of 0.7 V/, 2.5 eV and 14 fs, respectively. We note that for this value of the pulse duration, the field amplitude decays to zero at about 50 fs with a peak position at 25 fs. Figure 7 displays the time evolution of charges for the molecule and particle separately. We note that the total charge is preserved (here charge neutral complex) at all times, however, charge transfer can occur across the interface between the molecule and particle during the dynamics. As seen in Fig.7, the initial charge (before interacting with a laser pulse) of the molecule is slightly negative, reflecting ground state charge transfer to the molecule. The charges of the two subsystems oscillate with the laser frequency during interaction with the laser pulse. There is an effective transfer of electrons from the molecule to the particle during the early time of the dynamics (Figure 6), and the charge of the particle drops to about +0.15e at around 18 fs. After that the electron transfer changes direction and an effective transfer of electrons from the particle to the molecule is observed. At long times (after 50 fs), the charge on the molecule oscillates around a value of -0.75e. Figure 6 displays the negative charge on the molecule at long times (after 50 fs). The transfer of electrons and the formation of negative molecular ions for fullerene-based molecules have been experimentally demonstrated using SERS measurements, where the fullerene pentagonal pinch mode was found to change frequencies from that for the neutral fullerene to that for the fullerene anion after excitation of the gold/fullerene complex at 532 nm.SERS-PCBM-Schatz-2019 (33) We note that since the fullerene is still in contact with the gold nanoparticle at the end of the calculation the amount of charge residing on the molecule is not necessarily a full electron.

### v3

@@ Abstract
This study examines the Raman scattering and charge transfer properties of molecules adsorbed on the surface of a tetrahedral Au120 nanoparticle based on the time-dependent density functional tight-binding (TD-DFTB) method.
@@
Surface-enhanced Raman scattering (SERS) is a widely used phenomenon among various research communities where the scattering intensity of molecules can be increased by several orders of magnitude when they are adsorbed on the surface of plasmonic nanoparticles.SERS-VanDuyne-1997; SERS-Creighton-1977; SERS-VanDuyne-2008
The enhancement factors are as large as and therefore SERS can be studied on the single-molecule level.SERS-Emory-1997; SERS-Feld-1997; SERS-Schatz-2008; SERS-Kneipp-2008; SERS-CHeng-2019
The unexpected enhancement in the Raman signal from pyridine adsorbed on an electrochemically roughened silver electrode was the first evidence of SERS,SERS-McQuillan-1974 and this was attributed to the extensive surface coverage of molecules on the rough nanostructured surface, but later it was identified as a plasmon-based phenomenon.SERS-VanDuyne-1997; SERS-Creighton-1977
SERS has been successfully applied to several fields including biosensingBiosensing-VanDuyne-2003; Biosensing-Zhao-2008, drug analysis,Drug-Schatz-2009 and electrochemistryElectrochemical-Zen-2015; Electrochemical-VanDuyne-2015.
@@
Localized surface plasmon resonance (LSPR) excitation of the conduction electrons in metal nanoparticles of Au, Ag, and Cu with dimensions of 5-200 nanometers generates an enhanced electromagnetic field at the interface of the particle.LSPR-VanDuyne-2007
This enhanced field can modulate several optical processes at the surface including photocurrent,Photocurrent-Schatz-2009 luminescence,Exciton-Achermann-2010 and Raman scattering.SERS-VanDuyne-1997; SERS-Creighton-1977; SERS-VanDuyne-2008
Raman scattering scales with the local field as , where the numerator is the localized electromagnetic field and the denominator is the incident field.SERS-VanDuyne-2008
SERS enhancement factors due to electromagnetic effects can reach values up to .SERS-Schatz-2008
@@
Charge transfer (CT) excitations at optical frequency can also contribute to enhancements through a resonance Raman mechanism.Electronic-Schatz-2008
Enhancements due to static chemical effects can be up to a factor of 3-20 as demonstrated theoretically for pyridine adsorbed on the surface of a small tetrahedral Au particle.TDDFT-Schatz-2006.
Resonance CT can lead to larger enhancements, but usually the resonance CT and LSPR frequencies are sufficiently different, or the resonance CT transition is sufficiently broad, that one only has significant SERS enhancement from plasmon excitation and static CT.
@@
Previous theoretical works have addressed SERS for pyridine adsorbed on the surface of Ag and Au particles with enhancement factors up to based on the time-dependent density functional theory (TD-DFT).TDDFT-Schatz-2006; TDDFT-Schatz-2006
@@
Semi-empirical methods based on INDO have been demonstrated successfully for the efficient simulation of SERS for Ag-pyridine complexes.HyperRaman-Schatz-1992; INDO-Schatz-2017
As an alternative to the INDO calculations, the density functional tight-binding (DFTB) method may also provide efficient solutions for larger complexes.DFTB-Aradi-2020
@@
Two recent papers provide details of the DFTB method.DFTB-Aradi-2020; RTDFTB-Sanchez-2020
@@
Previous works have demonstrated the success of the DFTB method in various applications including biomolecules,DFTB-Bio-Wong-2019 metals,DFTB-Metal-2019; DFTB-Plasmon-Lucas-2020; DFTB-Plasmon-Sala-2022 semiconductors,DFTB-SemiConductor plasmonic photocatalysis,PDC-Sanchez-2022; PDC-Schatz-2023; PDC-Yu-2023, etc.
@@
The real-time time-dependent density functional tight-binding (RT-TD-DFTB) method allows us to study plasmon-driven charge and energy transfer processes at the metal-molecule interface by simulating laser-induced electron dynamics.RTDFTB-Sanchez-2020; PDC-Sanchez-2022; PDC-Schatz-2023
@@
This includes several processes like molecular bond dissociation and formation,PDC-Schatz-2023; PDC-Yu-2023 charge transfer (CT),Hot-Carrier-Erhart-2022 energy transfer,PDC-Sanchez-2022 spectral shift,SERS-PCBM-Schatz-2019 etc.
In the following, we study CT processes between Au120 and a fullerene (C60) after exciting plasmons at 2.5 eV with an external laser pulse.
This work was inspired by a recent study from the Van Duyne group, where the fullerene-based molecule [6,6]-phenyl-C61-butyric acid methyl ester (PCBM) was found to form negative ions in the presence of gold dimers when irradiated at 532 nm as revealed by SERS measurements.
@@
The paper is divided into two parts, in the first part, we explore Raman scattering for Au120-pyridine complexes and the second part contains plasmon-driven CT in a Au120-C60 complex.
@@ 2 Computational Details
To elucidate the accuracy of this method for predicting Raman spectra, we compare DFTB results with conventional results obtained with the B3LYP/TZVP level of DFT theory.
@@
To obtain a smooth Raman spectrum as a function of Raman shifts, the scattering activity is broadened by a Lorentzian function with a cm-1 width.
In real-time TD-DFTB (RT-TD-DFTB) the one-body reduced density matrix is propagated according to the Liouville-von Neumann equation in a nonorthogonal basisRTDFTB-Sanchez-2020
@@
with field amplitude , center of the pulse , pulse duration , and frequency . For calculating spectra, the density matrix evolves in time due to purely electronic dynamics where the nuclei are “clamped” in space. This treatment has been extensively used to study optical processes of material and molecules induced by external time-dependent electric fields.RTDFTB-Sanchez-2016; RTDFTB-2-Sanchez-2016; RTDFTB-Sanchez-2020 The Mulliken population is calculated as
@@
with an initial permanent dipole , and speed of light .
@@ 3 Results and Discussion
Au120-Pyridine Complexes:
We use a tetrahedral Au120 nanoparticle for all of these studies, which is a similar structure (but larger) to what has been considered in earlier SERS studies.TDDFT-Schatz-2006; TDDFT-Schatz-2006 This allows two distinct binding sites for an adsorbing pyridine molecule, either on one of the tips (V-complex) or in the middle of the triangle faces (S-complex).
@@
We optimize all geometries with a maximum force tolerance of 10-5 a.u. for the movement of atoms.
We attach pyridine on the tip (for V-complex) and surface (for S-complex) gold atoms of a previously optimized Au120 and further optimize, freezing the atoms of the particle.
@@
A smaller distance between the particle and molecule was also observed in a Au20-pyridine V-complex with TD-DFT calculations.TDDFT-Schatz-2006
@@
Furthermore, we note that an electronic temperature of 300K is used throughout the study.
@@
These spectral features are consistent with previous studies based on INDO/SCI and TD-DFT methods.INDO-Schatz-2017
@@
Absorption spectra for Au120 and Au120-pyridine complexes are shown in Figure 2(b).
The spectrum for the gold particle shows multiple peaks that are comparable to the spectrum of a tetrahedral Au20 investigated using the TD-DFT method.TDDFT-Schatz-2006
@@
However, the overall spectral intensity is increased with the particle size due to the large density of states available in Au120.
@@
We note that although a gold particle involving 120 atoms is large compared to the previous study with Au20, it is still too small sustain significant LSPR excitation.
@@
The spectrum for the S-complex is very similar to the isolated gold particle spectrum, whereas for the V-complex the absorption intensity is slightly increased due to CT excitations as is consistent with previous TD-DFT results.TDDFT-Schatz-2006
@@
We have used DFTB to compute the SERS enhancement for pyridine using low frequency (nominally static) calculations to determine the chemical enhancement factor, and with calculations at the plasmon energy of 2.5 eV for Au120-pyridine complexes to determine the electromagnetic enhancement factor.
We find that resonance CT excitations for the Au120-pyridine complexes involve small transition strengths that overlap with the dominant -band excitations and therefore can have negligible contributions to SERS.
@@
The Raman active modes from the TD-DFT calculation include a pair of ring deformation modes at 620 and 672 cm-1, a pair of ring breathing modes at 997 and 1039 cm-1, a symmetric C-H wag mode at 1242 cm-1, and a pair of ring stretching modes at 1604 and 1605 cm-1.
The two noticeable discrepancies between spectra are (i) unphysical modes that appear in the TD-DFTB results near 794 cm-1 and (ii) for the lower energy ring breathing mode (that appears at 941 cm-1 in the TD-DFTB result) the intensity is significantly suppressed compared to the others.
As will be discussed later the unphysical modes near 794 cm-1 are not enhanced in SERS, however, the lower energy ring breathing mode can be significantly enhanced even though it is vanishingly small for the isolated gas phase molecule (TD-DFTB result).
The normal Raman scattering spectra of pyridine adsorbed on the surface of Au120 are shown in Figure 4.
@@
These enhancement factors are comparable with the reported values (a factor 3 for the S-complex and 20 for the V-complex) for a Au20-pyridine complex based on earlier TD-DFT calculations.TDDFT-Schatz-2006
@@
Interestingly the lower energy ring breathing mode (944 cm-1) that has a small intensity for the isolated molecule (Fig.3) is now enhanced by a factor of 15 for the V-complex, while the higher energy ring breathing mode (1056 cm-1) shows a weak enhancement.
@@
Again, this trend in enhancements is consistent with the TD-DFT calculation as observed for Au20-pyridine complexes.TDDFT-Schatz-2006.
Further, the unphysical modes that appear in the isolated molecule near 794 cm-1 have not been enhanced for either complex.
Another mode with much higher enhancements is near 1459 cm-1. This peak was also seen in the TD-DFT work, and it appears in experimental SERS measurements on gold but not silver.TDDFT-Schatz-2006
The mode frequencies show blue-shifts for both complexes as observed in the TD-DFT calculations, however, the shifts are relatively smaller compared to the TD-DFT results.TDDFT-Schatz-2006
@@
Similar enhancement factors were also observed for Au20-pyridine complexes with the TD-DFT method.TDDFT-Schatz-2006 For the V-complex, the 941 and 1635 cm-1 modes are enhanced significantly, by a factor of and , respectively, while for the S-complex multiple modes are enhanced, with the maximum enhancement for 1636 cm-1 mode.
Further, a new Raman active mode appears in the S-complex at 1459 cm-1.
@@
Au120C60 Complex:
To generate a suitable geometry to study plasmon-driven charge transfer, we place a fullerene (C60) molecule on one of the triangular faces (as in the S-complex) of Au120 (previously optimized) and then the geometry is optimized freezing atoms of the particle.
The optimized geometry of Au120C60 is shown in Figure 6.
@@
Absorption spectra for C60 and Au120C60 are displayed in Figures 2 (a) and (b), respectively.
@@
These calculated peaks well represent the experimental peaks found at 3.8, 4.9, and 6 eV, respectively for the gas phase C60 molecule.C60-Shobatake-1996
The spectrum of the Au120C60 complex is dominated by the gold particle with two major peaks near 3.27 and 4.3 eV.
@@
Previous studies demonstrated a shift in LSPR peaks due to the interaction between plasmon and molecular resonances.Plasmon-Molecule-VanDuyne-2006; Plasmon-Molecule-VanDuyne-2007
@@
A linearly polarized (polarized along the Z-axis) laser pulse of the form given in Eq.2 is used to drive the system. Both molecule and particle couple with the external field. We use a driving field amplitude , frequency , and pulse duration of 0.7 V/, 2.5 eV and 14 fs, respectively. We note that for this value of the pulse duration, the field amplitude decays to zero at about 50 fs with a peak position at 25 fs. Figure 7 displays the time evolution of charges for the molecule and particle separately. We note that the total charge is preserved (here charge neutral complex) at all times, however, charge transfer can occur across the interface between the molecule and particle during the dynamics. As seen in Fig.7, the initial charge (before interacting with a laser pulse) of the molecule is slightly negative, reflecting ground state charge transfer to the molecule. The charges of the two subsystems oscillate with the laser frequency during interaction with the laser pulse. There is an effective transfer of electrons from the molecule to the particle during the early time of the dynamics (Figure 6), and the charge of the particle drops to about +0.15e at around 18 fs. After that the electron transfer changes direction and an effective transfer of electrons from the particle to the molecule is observed. At long times (after 50 fs), the charge on the molecule oscillates around a value of -0.75e. Figure 6 displays the negative charge on the molecule at long times (after 50 fs). The transfer of electrons and the formation of negative molecular ions for fullerene-based molecules have been experimentally demonstrated using SERS measurements, where the fullerene pentagonal pinch mode was found to change frequencies from that for the neutral fullerene to that for the fullerene anion after excitation of the gold/fullerene complex at 532 nm.SERS-PCBM-Schatz-2019 We note that since the fullerene is still in contact with the gold nanoparticle at the end of the calculation the amount of charge residing on the molecule is not necessarily a full electron.